# Continual Learning on CLIP via Incremental Prompt Tuning with Intrinsic Textual Anchors

### v1

@@ 4.4 Ablation Studies
Most importantly, for CIL, during the early incremental stages, models initially encounter only a tiny fraction of the total knowledge to be learned. The model has no idea what it would encounter in the future, thus learning current data would not take future data into consideration. However, when learning new data, some embeddins space is indeed needed to accommodate upcoming tasks (as illustrated in Fig. 1(a)), thus would cause representation drift (as shown in Fig. LABEL:fig:drift) of previously learned data. The use of textual prototypical contrastive loss helps mitigate this issue by maintaining more stable and meaningful embeddings throughout the incremental learning process.
@@
Prompt formulation. In Fig. LABEL:fig:ablate_prompt, we ablate the design of visual and textual prompts utilized in our proposed methods. Overall, both TPPT-V and TPPT-VT show improved performance with deeper prompts, where prompts are embedded into more advanced layers of Transformer blocks. Regarding the length of prompts, extended visual prompts benefits the prediction accuracy for both methods, while the performance of these methods appears relatively insensitive to the length of textual prompts.
@@ 5 Conclusion
Our work introduces TPPT-V and TPPT-VT, which leverages textual prototypes to guide CLIP visual prompts, further aligns text and vision via learned textual prompts and diversity regularization, respectively. Extensive experiments show that our simple, efficient design substantially reduces forgetting while improving adaptation to new tasks, representing a notable advance in CL.
@@
Acknowledgments and Disclosure of Funding
@@ Appendix A More Details of the Proposed Method
We illustrate the algorithm details of TPPT-VT in Algorithm 1. The training pipeline of TPPT-V follows the same structure as TPPT-VT, except that the textual features are obtained with frozen text encoder without textual prompts, and the training objectives for TPPT-V are detailed in Eq. (4).
@@
                   Aggregate the visual prompts ${\bm{p}}_{v}$ for given ${\mathbf{x}}$ via Eq. (2)
@@
                   Calculate per sample loss via Eq. (6)
@@ Appendix B Additional Experimental Details
We provide additional detials of our experiments in this appendix section. Reproducibility Statement. The source code will be made publicly available upon acceptance of the paper. We report experiment results averaged over 3 independent runs for reproducibility.
@@ B.1 Datasets
For a thorough evaluation of our methodologies, we compare the performance with previous methods across a variety of datasets. This includes three popular CIL datasets [54, 68, 67, 55]: CIFAR100 [32], ImageNet-R [10], and CUB200 [60], as well as two widely used datasets used for prompt tuning research [83, 73, 26, 27]: FGVCAircraft [47] and Stanford Cars [31]. Each dataset is splited into 10 tasks to formulate a CIL setup. We hereby provide a detailed description of these datasets.
CIFAR100 consists of 60,000 32$\times$32 color images in 100 classes, with 600 images per class. There are 500 training images and 100 testing images per class. Given that classes across different tasks may belong to the same superclass, there is a potential overlap in similarities between tasks.
@@ B.2 Compared Methods
In our experiments, we conduct a comprehensive comparison of our methods against a variety of previous methods. Firstly, we assess the zero-shot performance of the pre-trained CLIP model (ZS-CLIP) [52]. This is followed by comparisons with two widely recognized prompt tuning methods: CoOp [83] and VPT-CL (our implementation of VPT [24]), adapted to CIL senario. Further, we compare against several popular CL methods: L2P [68], DualPrompt [67], CODA-P [55], AttriCLIP [64], ZSCL [77], MoE-Adapter [74], InfLoRA [38], and PROOF [81], to provide a comprehensive evaluation of our methods.
ZS-CLIP. We use the pre-trained frozen image encoder and text encoder to obtain visual features and textual features, and use textual features as a classifier (Eq. (1)) to make predictions. Specifically, textual features are obtained through hand-crafted prompt templates as suggested by OpenAI  [^2] https://github.com/openai/CLIP with class names.
CoOp. CoOp introduces a method of prepending learnable prompt tokens to the text encoder of the CLIP model. This strategy is designed to adapt the pre-trained model for various downstream tasks. In our experiments, we adapt the CoOp design to fit the CIL scenario, that continually trains the textual prompts over a sequence of tasks.
VPT-CL. VPT is a popular prompt tuning method designed specifically for Vision Transformer (ViT) model. In our study, we adapt VPT for use with the pre-trained CLIP model in a CIL context, we name it VPT-CL. Specifically, we discard the classification head of VPT and make predictions using textual features as in ZS-CLIP, and we adapt this implementation to CIL scenario that continually train the visual prompts.
@@
AttriCLIP. AttriCLIP enhances the textual branch of the CLIP model through the learning of text attribute prompts. These prompts are chosen from the proposed attribute bank, to encode boarder knowledge of textual information.
@@
Table 6: Average and final accuracy of different methods with different buffer sizes. All methods are trained with the same pre-trained weights, and employing the CLIP model with a ViT-B/16 backbone. $\bar{\mathcal{A}}$ denotes the average of test accuracy over incremental stages, $\mathcal{A}_{T}$ denotes the test accuracy at the last incremental stage. The best and secondary performances are indicated in red and blue, respectively.
@@
Figure 15: Comparisons to AttriCLIP on CIFAR-100 using ViT-L/14 with pre-trained weight from OpenAI[52].
@@
Table 9: Average and final accuracy of different methods using OpenAI [52] pre-trained weight. All methods are trained with identical exemplar sizes, utilizing the same pre-trained weights, and employing the CLIP model with a ViT-B/16 backbone. $\bar{\mathcal{A}}$ denotes the average of test accuracy over incremental stages, $\mathcal{A}_{T}$ denotes the test accuracy at the last incremental stage. The best and secondary performances are indicated in red and blue, respectively.
@@ Appendix C More Experimental Results and Ablation Studies
Additional comparison with AttriCLIP [64]. For a fair comparison, we compare our methods with AttriCLIP reported results under the same experiment settings [64] on CIFAR-100 in Fig. 15. The backbone model employed was CLIP with a ViT-L/14 architecture, utilizing pre-trained weights from OpenAI. Throughout the incremental learning process, our approach demonstrated superior performance, consistently outperforming AttriCLIP.
@@
Experiment results on Upper-bound performance. In complement to the incremental learning performance in Table 1 and Fig. LABEL:fig:exp_task, we further report the Upper-bound performance of our methods in Table 7, measuring the learning capacity of the model under non-incremental setting. Notably, for fine-grained datasets such as CUB, Aircraft, and Cars, the performance gap between the final accuracy reported in Table 1 and the Upper-bound performance is marginal, highlights the effectiveness of the introduction of textual prototypes, which greatly stabilize the training during the incremental phases.
@@
Studies on Forgetting. In Table 1, we present the average and final accuracies of our proposed methods, alongside comparisons with previous works. Additionally, we evaluate the average forgetting of our methods in Table 5 following [68, 67, 55]. We formally denote the forgetting metric for a given task as: $\mathcal{F}_{t}=\frac{1}{t-1}\sum_{\tau=1}^{t-1}\max_{\tau^{\prime}\in\{1,\cdots,t-1\}}(\mathcal{A}_{\tau^{\prime},\tau}-\mathcal{A}_{t,\tau}),$ where $\mathcal{A}_{t,\tau}$ represents the accuracy on task $\tau$ after training on task $t$. We then evaluate average of forgetting over tasks ~$\tilde{\mathcal{F}}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{F}_{t}$. The results demonstrate that our methods not only consistently outperform the previous approaches but also exhibit low levels of forgetting, thus effectively preventing catastrophic forgetting. Benefited from our design of learning with textual prototypes, our method is inherently less prone to forget. Enriching textual prototypes with textual diversity not only enhances the stability and efficacy of the training process but also slightly improves the forgetting measure. While CoOp achieves lower forgetting on the Aircraft dataset, it is important to note the substantial difference in prediction accuracy favoring our methods. Given this considerable performance gap, our proposed methods stand out as superior in terms of overall efficacy.
@@
Analyses on reply buffer. To provide a comprehensive understanding, we explore the impact of different buffer sizes on performance. These additional evaluations, including comparisons with previous works, are detailed in Table 6. The results from these experiments indicate that our proposed methods consistently outperform or are comparable to previous methods, demonstrating robustness irrespective of the chosen buffer size. A notable aspect of our methods is the effectiveness even with a limited number of exemplars. For instance, when using only 5 exemplars per task, TPPT-VT achieves an average accuracy of 79.29 on the CUB dataset. This performance is on par with that of previous works using larger number of exemplars, which attain an average accuracy of 79.69 using 15 exemplars per task. Such results show the efficiency and competitiveness of our methods even under constrained exemplar conditions.
@@
Experiment results with larger number of tasks. Evaluation results in Table 1 and Fig. LABEL:fig:exp_task were conducted under our default experimental setup, where each dataset is divided into 10 tasks for incremental learning. In Fig 14, we extend our evaluation to a more challenging experiment setting by splitting each dataset into 20 tasks, which increases the complexity of the learning process and makes the model more easy to forget. Despite the increased complexity, our methods, which leverage learning with textual prototypes and the diversity regularization of these prototypes, consistently outperform previous methods in the 20-task split setting.
@@
Experiment results with varying class orders. In Table 8, we perform 5 independent runs and report the average final accuracy and standard deviation. Our methods is robust to different class orderings, showing its robust performance across varied learning scenarios.
@@
Computational efficiency of our methods. In addition to the discussion of our parameters in Sec. 4.4, we hereby discuss the computation efficiency of our methods. All experiments are run using one NVIDIA RTX3090 GPU. CODA-P uses similar design of incremental prompts like us, takes approximately 8ms for training and 5ms for inference per image. TPPT-V takes a comparable or less amount of time, as it does not require additional learnable attention vectors, yet it delivers a significant performance boost of 11% in final accuracy averaged across all datasets compared to CODA-P. TPPT-VT additionally learns textual prompts, and further increases final accuracy on the Aircraft dataset by 13%, with only an additional 1.5ms required for both training and inference compared to TPPT-V. While these methods train prompt tokens across multiple layers of the CLIP model, PROOF employs a different approach by training a set of incremental projectors appended to the CLIP encoders, resulting in less processing time. Our methods surpass PROOF by anchoring previous knowledge with textual prototypes and regularizing the learning of multi-modal prompts, thereby enhancing performance without significantly increasing computational overhead.
@@
Experiment results using OpenAI pre-trained CLIP weight. In our main paper, we mainly report the performance using pre-trained weight from LAION-400M [21]. To demonstrate the robustness of our proposed methods, we further evaluate our performance using pre-trained weight from OpenAI [52] in Table 9, and we illustrate the test accuracy over incremental stages in Fig. LABEL:supp:openai_exp_fig. Our experiment results demonstrate that our proposed methods are insensitive to the selection of pre-trained weight, consistently outperforms previous methods.

### v2

@@ 4.4 Ablation Studies
Most importantly, for CIL, during the early incremental stages, models initially encounter only a tiny fraction of the total knowledge to be learned. The model has no idea what it will encounter in the future; thus, learning current data does not take future data into consideration. However, when learning new data, some embedding space is indeed needed to accommodate upcoming tasks (as illustrated in Fig. 1(a)), which causes representation drift (as shown in Fig. LABEL:fig:drift) of previously learned data. The use of textual prototypical contrastive loss helps mitigate this issue by maintaining more stable and meaningful embeddings throughout the incremental learning process.
@@
Prompt formulation. In Fig. LABEL:fig:ablate_prompt, we ablate the design of visual and textual prompts utilized in our proposed methods. Overall, both TPPT-V and TPPT-VT show improved performance with deeper prompts, where prompts are embedded into more advanced layers of Transformer blocks. Regarding the length of prompts, extended visual prompts benefit the prediction accuracy for both methods, while the performance of these methods appears relatively insensitive to the length of textual prompts.
@@ 5 Conclusion
Our work introduces TPPT-V and TPPT-VT, which leverage textual prototypes to guide CLIP visual prompts, further aligning text and vision via learned textual prompts and diversity regularization, respectively. Extensive experiments show that our simple, efficient design substantially reduces forgetting while improving adaptation to new tasks, representing a notable advance in CL.
@@
Broader Impact Statement
Acknowledgments
@@ Appendix A More Details of the Proposed Method
We illustrate the algorithm details of TPPT-VT in Algorithm 1. The training pipeline of TPPT-V follows the same structure as TPPT-VT, except that the textual features are obtained with the frozen text encoder without textual prompts, and the training objectives for TPPT-V are detailed in Equation 4.
@@
         Aggregate the visual prompts ${\bm{p}}_{v}$ for given $\mathbf{x}$} via Equation 2
@@
         Calculate per sample loss via Equation 6
@@ Appendix B Additional Experimental Details
We provide additional details of our experiments in this appendix section. Reproducibility Statement. The source code will be made publicly available upon acceptance of the paper.
@@ B.1 Datasets
For a thorough evaluation of our methodologies, we compare the performance with previous methods across a variety of datasets. This includes three popular CIL datasets (Rebuffi et al., 2017; Wang et al., 2022e; d; Smith et al., 2023): CIFAR100 (Krizhevsky et al., 2009), ImageNet-R (Deng et al., 2009), TinyImageNet (Le & Yang, 2015), and CUB200 (Wah et al., 2011), as well as two widely used datasets used for prompt tuning research (Zhou et al., 2022c; Yao et al., 2023; Khattak et al., 2023a; b): FGVCAircraft (Maji et al., 2013) and Stanford Cars (Krause et al., 2013). Each dataset is split into 10 tasks to formulate a CIL setup. We hereby provide a detailed description of these datasets.
CIFAR100 consists of 60,000 $\times32$ color images in 100 classes, with 600 images per class. There are 500 training images and 100 testing images per class. Given that classes across different tasks may belong to the same superclass, there is a potential overlap in similarities between tasks.
@@
TinyImageNet is a downsampled subset of ImageNet containing 200 classes, each resized to $64\times 64$ pixels. It is widely used as a challenging benchmark for evaluating continual learning algorithms due to its larger class set and greater intra-class variability compared to CIFAR100.
@@ B.2 Compared Methods
In our experiments, we conduct a comprehensive comparison of our methods against a variety of previous methods. Firstly, we assess the zero-shot performance of the pre-trained CLIP model (ZS-CLIP) (Radford et al., 2021). This is followed by comparisons with two widely recognized prompt tuning methods: CoOp (Zhou et al., 2022c) and VPT-CL (our implementation of VPT (Jia et al., 2022)), adapted to the CIL scenario. Further, we compare against several popular CL methods: L2P (Wang et al., 2022e), DualPrompt (Wang et al., 2022d), CODA-P (Smith et al., 2023), AttriCLIP (Wang et al., 2023), ZSCL (Zheng et al., 2023), MoE-Adapter (Yu et al., 2024), InfLoRA (Liang & Li, 2024), and PROOF (Zhou et al., 2023c), to provide a comprehensive evaluation of our methods.
ZS-CLIP. We use the pre-trained frozen image encoder and text encoder to obtain visual features and textual features, and use textual features as a classifier (Equation 1) to make predictions. Specifically, textual features are obtained through hand-crafted prompt templates as suggested by OpenAI  [^2] https://github.com/openai/CLIP with class names.
CoOp. CoOp introduces a method of prepending learnable prompt tokens to the text encoder of the CLIP model. This strategy is designed to adapt the pre-trained model for various downstream tasks. In our experiments, we adapt the CoOp design to fit the CIL scenario, which continually trains the textual prompts over a sequence of tasks.
VPT-CL. VPT is a popular prompt tuning method designed specifically for the Vision Transformer (ViT) model. In our study, we adapt VPT for use with the pre-trained CLIP model in a CIL context, and we name it VPT-CL. Specifically, we discard the classification head of VPT and make predictions using textual features as in ZS-CLIP, and we adapt this implementation to the CIL scenario that continually trains the visual prompts.
@@
AttriCLIP. AttriCLIP enhances the textual branch of the CLIP model through the learning of text attribute prompts. These prompts are chosen from the proposed attribute bank to encode broader knowledge of textual information.
@@
Table 6: Average and final accuracy of different methods with different buffer sizes. All methods are trained with the same pre-trained weights and employ the CLIP model with a ViT-B/16 backbone. $\bar{\mathcal{A}}$ denotes the average of test accuracy over incremental stages, $\mathcal{A}_{T}$ denotes the test accuracy at the last incremental stage. The best and secondary performances are indicated in red and blue, respectively.
@@
Figure 21: Comparisons to AttriCLIP on CIFAR-100 using ViT-L/14 with pre-trained weight from OpenAI(Radford et al., 2021).
@@
Table 9: Average and final accuracy of different methods using OpenAI (Radford et al., 2021) pre-trained weight. All methods are trained with identical exemplar sizes, utilizing the same pre-trained weights, and employing the CLIP model with a ViT-B/16 backbone. $\bar{\mathcal{A}}$ denotes the average of test accuracy over incremental stages, $\mathcal{A}_{T}$ denotes the test accuracy at the last incremental stage. The best and secondary performances are indicated in red and blue, respectively.
@@ Appendix C More Experimental Results and Ablation Studies
Additional comparison with AttriCLIP (Wang et al., 2023). For a fair comparison, we compare our methods with AttriCLIP reported results under the same experiment settings (Wang et al., 2023) on CIFAR-100 in Fig. 21. The backbone model employed was CLIP with a ViT-L/14 architecture, utilizing pre-trained weights from OpenAI. Throughout the incremental learning process, our approach demonstrated superior performance, consistently outperforming AttriCLIP.
@@
Experiment results on Upper-bound performance. In complement to the incremental learning performance in Table 1 and Fig. LABEL:fig:exp_task, we further report the Upper-bound performance of our methods in Table 7, measuring the learning capacity of the model under a non-incremental setting. Notably, for fine-grained datasets such as CUB, Aircraft, and Cars, the performance gap between the final accuracy reported in Table 1 and the Upper-bound performance is marginal, highlighting the effectiveness of the introduction of textual prototypes, which greatly stabilizes the training during the incremental phases.
@@
Studies on Forgetting. In Table 1, we present the average and final accuracies of our proposed methods, alongside comparisons with previous works. Additionally, we evaluate the average forgetting of our methods in Table 5 following (Wang et al., 2022e; d; Smith et al., 2023). We formally denote the forgetting metric for a given task as: $\mathcal{F}_{t}=\frac{1}{t-1}\sum_{\tau=1}^{t-1}\max_{\tau^{\prime}\in\{1,\cdots,t-1\}}(\mathcal{A}_{\tau^{\prime},\tau}-\mathcal{A}_{t,\tau}),$ where $\mathcal{A}_{t,\tau}$ represents the accuracy on task $\tau$ after training on task $t$. We then evaluate average of forgetting over tasks $\tilde{\mathcal{F}}=\frac{1}{T}\sum_{t=1}^{T}\mathcal{F}_{t}$. The results demonstrate that our methods not only consistently outperform the previous approaches but also exhibit low levels of forgetting, thus effectively preventing catastrophic forgetting. Benefited from our design of learning with textual prototypes, our method is inherently less prone to forgetting. Enriching textual prototypes with textual diversity not only enhances the stability and efficacy of the training process but also slightly improves the forgetting measure. While CoOp achieves lower forgetting on the Aircraft dataset, it is important to note the substantial difference in prediction accuracy favoring our methods. Given this considerable performance gap, our proposed methods stand out as superior in terms of overall efficacy.
@@
Analyses on reply buffer. To provide a comprehensive understanding, we explore the impact of different buffer sizes on performance. These additional evaluations, including comparisons with previous works, are detailed in Table 6. The results from these experiments indicate that our proposed methods consistently outperform or are comparable to previous methods, demonstrating robustness irrespective of the chosen buffer size. A notable aspect of our methods is the effectiveness even with a limited number of exemplars. For instance, when using only 5 exemplars per task, TPPT-VT achieves an average accuracy of 79.29 on the CUB dataset. This performance is on par with that of previous works using a larger number of exemplars, which attain an average accuracy of 79.69 using 15 exemplars per task. Such results show the efficiency and competitiveness of our methods even under constrained exemplar conditions.
@@
Experiment results with a larger number of tasks. Evaluation results in Table 1 and Fig. LABEL:fig:exp_task were conducted under our default experimental setup, where each dataset is divided into 10 tasks for incremental learning. In Fig. 20, we extend our evaluation to a more challenging experiment setting by splitting each dataset into 20 tasks, which increases the complexity of the learning process and makes the model more prone to forgetting. Despite the increased complexity, our methods, which leverage learning with textual prototypes and the diversity regularization of these prototypes, consistently outperform previous methods in the 20-task split setting.
@@
Experiment results with varying class orders. In Table 8, we perform 5 independent runs and report the average final accuracy and standard deviation. Our method is robust to different class orderings, showing its robust performance across varied learning scenarios.
@@
Computational efficiency of our methods. In addition to the discussion of our parameters in Sec. 4.4, we hereby discuss the computation efficiency of our methods. All experiments are run using one NVIDIA RTX3090 GPU. CODA-P uses a similar design of incremental prompts to us, takes approximately 8ms for training and 5ms for inference per image. TPPT-V takes a comparable or less amount of time, as it does not require additional learnable attention vectors, yet it delivers a significant performance boost of 11% in final accuracy averaged across all datasets compared to CODA-P. TPPT-VT additionally learns textual prompts and further increases final accuracy on the Aircraft dataset by 13%, with only an additional 1.5ms required for both training and inference compared to TPPT-V. While these methods train prompt tokens across multiple layers of the CLIP model, PROOF employs a different approach by training a set of incremental projectors appended to the CLIP encoders, resulting in less processing time. Our methods surpass PROOF by anchoring previous knowledge with textual prototypes and regularizing the learning of multi-modal prompts, thereby enhancing performance without significantly increasing computational overhead.
@@
Experiment results using OpenAI pre-trained CLIP weight. In our main paper, we mainly report the performance using pre-trained weights from LAION-400M (Ilharco et al., 2021). To demonstrate the robustness of our proposed methods, we further evaluate our performance using pre-trained weights from OpenAI (Radford et al., 2021) in Table 9, and we illustrate the test accuracy over incremental stages in Fig. LABEL:supp:openai_exp_fig. Our experiment results demonstrate that our proposed methods are insensitive to the selection of pre-trained weights, consistently outperforming previous methods.